# Task-free Adaptive Meta Black-box Optimization

### v1

@@ 3.2 Meta-Strategy Architecture
ABOM implements a differentiable meta-strategy $\hat{\mathbf{P}}^{(t)}=\pi_{\bm{\theta}}(\mathbf{P}^{(t)},\mathbf{F}^{(t)})$ (Fig. 2, Bottom) that learns evolutionary operators via attention mechanisms Vaswani et al. (2017). At generation $t$, the population $\mathbf{P}^{(t)}=\left[\mathbf{p}_{1}^{(t)\top};\;\dots;\;\mathbf{p}_{N}^{(t)\top}\right]\in\mathbb{R}^{N\times d}$ represents $N$ candidate solutions in the search space, where the individual $\mathbf{p}_{i}^{(t)}\in\mathbb{R}^{d}$ is a solution vector. The fitness values $\mathbf{F}^{(t)}=\left[f_{T}(\mathbf{p}_{1}^{(t)});...;f_{T}(\mathbf{p}_{N}^{(t)})\right]\in\mathbb{R}^{N}$ are scalar evaluations $f_{T}(\mathbf{p}_{i}^{(t)})$ obtained via black-box queries to the target objective $f_{T}(\cdot)$, with lower values indicating better solutions. Given the population-fitness pair, ABOM generates offspring through three unified modules:
@@
where $\mathrm{MLP}_{\theta_{c}}(\mathbf{z})=\mathbf{W}_{2}\tanh(\mathbf{W}_{1}\mathbf{z}+\mathbf{b}_{1})+\mathbf{b}_{2}$ with $\mathbf{W}_{1}\in\mathbb{R}^{d\times d_{M}}$, $\mathbf{b}_{1}\in\mathbb{R}^{d_{M}}$, $\mathbf{W}_{2}\in\mathbb{R}^{d_{M}\times d}$, $\mathbf{b}_{2}\in\mathbb{R}^{d}$. Dropout with rate $p_{C}$ is applied to the hidden layer during both adaptive parameter learning and inference. The mechanism ensures persistent exploration through controlled randomness and is consistently maintained across all stochastic operations in ABOM. The term $\mathbf{A}^{(t)}\mathbf{P}^{(t)}$ computes an adaptive recombination pool: each row $\sum_{j=1}^{N}\mathbf{A}^{(t)}_{i,j}\mathbf{p}_{j}^{(t)}$ represents a context-aware blend of parent solutions, where weights $\mathbf{A}^{(t)}_{i,j}$ dynamically balance proximity in solution space and fitness-driven selection pressure.
@@
where $\mathbf{W}^{Q\!M},\mathbf{W}^{K\!M}\in\mathbb{R}^{1\times d_{A}}$, and $\mathrm{MLP}_{\theta_{m}}(\mathbf{z})=\mathbf{W}_{4}\tanh(\mathbf{W}_{3}\mathbf{z}+\mathbf{b}_{3})+\mathbf{b}_{4}$ with $\mathbf{W}_{3}\in\mathbb{R}^{d\times d_{M}}$, $\mathbf{b}_{3}\in\mathbb{R}^{d_{M}}$, $\mathbf{W}_{4}\in\mathbb{R}^{d_{M}\times d}$, $\mathbf{b}_{4}\in\mathbb{R}^{d}$. Following the same exploration principle as crossover, dropout with rate $p_{M}$ is applied during inference to maintain persistent exploration. The mutation matrix $\mathbf{M}^{(t)}\in\mathbb{R}^{d\times d}$ dynamically models gene-wise dependencies: each entry $\mathbf{M}^{(t)}_{j,k}$ quantifies the interaction strength between the $j$-th and $k$-th dimensions, enabling context-aware perturbations. Finally, offspring are concatenated as:
@@ 3.3 Adaptive Parameter Learning
As shown in Fig. 2 (Top), ABOM’s optimization loop comprises: 1) Initialization: The initial population $\mathbf{P}^{(0)}$ is randomly generated by Latin hypercube sampling; 2) Reproduction: Offspring $\hat{\mathbf{P}}^{(t)}$ are generated via $\hat{\mathbf{P}}^{(t)}=\pi_{\bm{\theta}}(\mathbf{P}^{(t)},\mathbf{F}^{(t)})$; 3) Evaluation: Fitness values $\hat{\mathbf{F}}^{(t)}$ are computed for $\hat{\mathbf{P}}^{(t)}$; 4) Elitism Deb et al. (2002): The elite archive $\mathbf{E}^{(t)}\in\mathbb{R}^{N\times d}$, formed by the top $N$ individuals from $\mathbf{P}^{(t)}\cup\hat{\mathbf{P}}^{(t)}$, and their fitness values $\mathbf{F}_{\mathbf{E}}^{(t)}\in\mathbb{R}^{N}$, are carried over to the next generation; 5) Parameter adaptation: $\bm{\theta}$ is updated via adaptive parameter learning. The pseudocode of ABOM can be found in the Appendix C (Alg. 1). Crucially, ABOM performs adaptive parameter learning by minimizing the distance between offspring and the elite archive:
@@ 3.4 Computational Complexity and Convergence Analysis
The computational cost of ABOM is primarily dominated by the selection, crossover, and mutation. The selection matrix (Eq. 5) incurs complexity $O(Ndd_{A}+N^{2}d_{A})$, where $N$ is the population size, $d$ the search space dimension, and $d_{A}$ the attention dimension. The MLP of the crossover (Eq. 6) contributes $O(Nd_{A}d_{M}+Nd_{M}d)$, with $d_{M}$ the hidden dimension of the MLP. The mutation (Eq. 7) contributes $O(d^{2}d_{A}+dd_{A}d_{M})$. Summing these, the total complexity is:
@@ Appendix D Convergence Analysis of ABOM
Define $f^{*}_{t}=\min_{\mathbf{x}\in\mathbf{E}^{(t)}}f_{T}(\mathbf{x})$, where $\mathbf{E}^{(t)}$ is the elite archive containing the top $N$ individuals from $\mathbf{P}^{(t)}\cup\hat{\mathbf{P}}^{(t)}$. Let $\mathcal{F}_{t}=\sigma(\mathbf{P}^{(0)},\dots,\mathbf{P}^{(t)},\theta^{(0)},\dots,\theta^{(t)})$ be the filtration representing all information up to generation $t$. The elitism mechanism ensures $f^{*}_{t+1}\leq f^{*}_{t}$ almost surely, implying $\mathbb{E}[f^{*}_{t+1}\mid\mathcal{F}_{t}]\leq f^{*}_{t}$. Since $f_{T}$ is bounded on the compact set $\mathcal{X}$, the sequence $\{f^{*}_{t},\mathcal{F}_{t}\}$ forms a lower-bounded supermartingale. By the martingale convergence theorem Hall and Heyde (2014), $f^{*}_{t}$ converges with probability 1 to the random variable $f^{*}_{\infty}\geq f^{*}$. To establish global convergence, we need to prove $f^{*}_{\infty}=f^{*}$ with probability 1.
@@
$\mathrm{MLP}_{\theta_{c}^{*}}(\mathbf{A}^{(t)}\mathbf{p}_{i}^{(t)})=\mathbf{W}_{2}\tanh(\mathbf{W}_{1}\mathbf{A}^{(t)}\mathbf{p}_{i}^{(t)}+\mathbf{b}_{1})+\mathbf{b}_{2}=\mathbf{v}.$ (16)
@@
$\mathrm{MLP}_{\theta_{m}^{*}}(\mathbf{M}_{i}^{(t)}\mathbf{p}_{i}^{{}^{\prime}(t)})=\mathbf{W}_{4}\tanh(\mathbf{W}_{3}\mathbf{M}_{i}^{(t)}\mathbf{p}_{i}^{{}^{\prime}(t)}+\mathbf{b}_{3})+\mathbf{b}_{4}=\mathbf{w}.$ (21)
@@
Let $\gamma=\mu_{c}\mu_{m}>0$, which is a positive constant independent of $t$ and depends only on hyperparameters $p_{C}$, $p_{M}$, $d_{M}$, and the adaptive parameter learning. Finally, for $N$ independent offspring:
@@ Appendix I Sensitivity Analysis of Learning Rate
We introduce ABOM-PT, a pre-trained variant of ABOM, where meta-optimization knowledge is distilled from the training set. Specifically, ABOM is executed on 10 source tasks for $T=250$ iterations per task, generating $2500$ prior training samples $\mathcal{M}=\{(P_{k}^{(t)},F_{k}^{(t)})\mid P_{k}^{(t)}\in\mathbb{R}^{N\times d},F_{k}^{(t)}\in\mathbb{R}^{N},k=1,...,K,t=1,...,T\}$. The pre-training objective minimizes the prediction error of population evolution:

### v2

@@ 3.2 Meta-Strategy Architecture
ABOM implements a differentiable meta-strategy $\hat{\mathbf{P}}^{(t)}=\pi_{\bm{\theta}}(\mathbf{P}^{(t)},\mathbf{F}^{(t)})$ (Fig. 2, Bottom) that learns evolutionary operators via attention mechanisms Vaswani et al. (2017). At generation $t$, the population $\mathbf{P}^{(t)}=\left[\mathbf{p}_{1}^{(t)\top};\;\dots;\;\mathbf{p}_{N}^{(t)\top}\right]\in\mathbb{R}^{N\times d}$ represents NN candidate solutions in the search space, where the individual $\mathbf{p}_{i}^{(t)}\in\mathbb{R}^{d}$ is a solution vector. The fitness values $\mathbf{F}^{(t)}=\left[f_{T}(\mathbf{p}_{1}^{(t)});...;f_{T}(\mathbf{p}_{N}^{(t)})\right]\in\mathbb{R}^{N}$ are scalar evaluations $f_{T}(\mathbf{p}_{i}^{(t)})$ obtained via black-box queries to the target objective $f_{T}(\cdot)$, with lower values indicating better solutions. Given the population-fitness pair, ABOM generates offspring through three unified modules:
@@
where $\mathrm{MLP}_{\theta_{c}}(\mathbf{z})=\tanh(\mathbf{z}\mathbf{W}_{1}+\mathbf{b}_{1})\mathbf{W}_{2}+\mathbf{b}_{2}$ with $\mathbf{W}_{1}\in\mathbb{R}^{d\times d_{M}}$, $\mathbf{b}_{1}\in\mathbb{R}^{d_{M}}$, $\mathbf{W}_{2}\in\mathbb{R}^{d_{M}\times d}$, $\mathbf{b}_{2}\in\mathbb{R}^{d}$. Dropout with rate $p_{C}$ is applied to the hidden layer during both adaptive parameter learning and inference. The mechanism ensures persistent exploration through controlled randomness and is consistently maintained across all stochastic operations in ABOM. The term $\mathbf{A}^{(t)}\mathbf{P}^{(t)}$ computes an adaptive recombination pool: each row $\sum_{j=1}^{N}\mathbf{A}^{(t)}_{i,j}\mathbf{p}_{j}^{(t)}$ represents a context-aware blend of parent solutions, where weights $\mathbf{A}^{(t)}_{i,j}$ dynamically balance proximity in solution space and fitness-driven selection pressure.
@@
where $\mathbf{W}^{Q\!M},\mathbf{W}^{K\!M}\in\mathbb{R}^{1\times d_{A}}$, and $\mathrm{MLP}_{\theta_{m}}(\mathbf{z})=\tanh(\mathbf{z}\mathbf{W}_{3}+\mathbf{b}_{3})\mathbf{W}_{4}+\mathbf{b}_{4}$ with $\mathbf{W}_{3}\in\mathbb{R}^{1\times d_{M}}$, $\mathbf{b}_{3}\in\mathbb{R}^{d_{M}}$, $\mathbf{W}_{4}\in\mathbb{R}^{d_{M}}$, $\mathbf{b}_{4}\in\mathbb{R}$. Following the same exploration principle as crossover, dropout with rate $p_{M}$ is applied during inference to maintain persistent exploration. The mutation matrix $\mathbf{M}^{(t)}\in\mathbb{R}^{d\times d}$ dynamically models gene-wise dependencies: each entry $\mathbf{M}^{(t)}_{j,k}$ quantifies the interaction strength between the $j$-th and $k$-th dimensions, enabling context-aware perturbations. Finally, offspring are concatenated as:
@@ 3.3 Adaptive Parameter Learning
As shown in Fig. 2 (Top), ABOM’s optimization loop comprises: 1) Initialization: The initial population $\mathbf{P}^{(0)}$ is randomly generated by Latin hypercube sampling; 2) Reproduction: Offspring $\hat{\mathbf{P}}^{(t)}$ are generated via $\hat{\mathbf{P}}^{(t)}=\pi_{\bm{\theta}}(\mathbf{P}^{(t)},\mathbf{F}^{(t)})$; 3) Evaluation: Fitness values $\hat{\mathbf{F}}^{(t)}$ are computed for $\hat{\mathbf{P}}^{(t)}$; 4) Elitism Deb et al. (2002): The elite archive $\mathbf{E}^{(t)}\in\mathbb{R}^{N\times d}$, formed by the top NN individuals from $\mathbf{P}^{(t)}\cup\hat{\mathbf{P}}^{(t)}$, and their fitness values $\mathbf{F}_{\mathbf{E}}^{(t)}\in\mathbb{R}^{N}$, are carried over to the next generation; 5) Parameter adaptation: $\bm{\theta}$ is updated via adaptive parameter learning. The pseudocode of ABOM can be found in the Appendix C (Alg. 1). Crucially, ABOM performs adaptive parameter learning by minimizing the distance between offspring and the elite archive:
@@ 3.4 Computational Complexity and Convergence Analysis
The computational cost of ABOM is primarily dominated by the selection, crossover, and mutation. The selection matrix (Eq. 5) incurs complexity $O(Ndd_{A}+N^{2}d_{A})$, where NN is the population size, $d$ the search space dimension, and $d_{A}$ the attention dimension. The MLP of the crossover (Eq. 6) contributes $O(Nd_{A}d_{M}+Nd_{M}d)$, with $d_{M}$ the hidden dimension of the MLP. The mutation (Eq. 7) contributes $O(d^{2}d_{A}+dd_{A}d_{M})$. Summing these, the total complexity is:
@@ Appendix D Convergence Analysis of ABOM
Define $f^{*}_{t}=\min_{\mathbf{x}\in\mathbf{E}^{(t)}}f_{T}(\mathbf{x})$, where $\mathbf{E}^{(t)}$ is the elite archive containing the top NN individuals from $\mathbf{P}^{(t)}\cup\hat{\mathbf{P}}^{(t)}$. Let $\mathcal{F}_{t}=\sigma(\mathbf{P}^{(0)},\dots,\mathbf{P}^{(t)},\theta^{(0)},\dots,\theta^{(t)})$ be the filtration representing all information up to generation $t$. The elitism mechanism ensures $f^{*}_{t+1}\leq f^{*}_{t}$ almost surely, implying $\mathbb{E}[f^{*}_{t+1}\mid\mathcal{F}_{t}]\leq f^{*}_{t}$. Since $f_{T}$ is bounded on the compact set $\mathcal{X}$, the sequence $\{f^{*}_{t},\mathcal{F}_{t}\}$ forms a lower-bounded supermartingale. By the martingale convergence theorem Hall and Heyde (2014), $f^{*}_{t}$ converges with probability 1 to the random variable $f^{*}_{\infty}\geq f^{*}$. To establish global convergence, we need to prove $f^{*}_{\infty}=f^{*}$ with probability 1.
@@
$\mathrm{MLP}_{\theta_{c}^{*}}(\mathbf{A}^{(t)}\mathbf{p}_{i}^{(t)})=\tanh(\mathbf{A}^{(t)}\mathbf{p}_{i}^{(t)}\mathbf{W}_{1}+\mathbf{b}_{1})\mathbf{W}_{2}+\mathbf{b}_{2}=\mathbf{v}.$ (16)
@@
$\mathrm{MLP}_{\theta_{m}^{*}}(\mathbf{M}_{i}^{(t)}\mathbf{p}_{i}^{{}^{\prime}(t)})=\tanh(\mathbf{M}_{i}^{(t)}\mathbf{p}_{i}^{{}^{\prime}(t)}\mathbf{W}_{3}+\mathbf{b}_{3})\mathbf{W}_{4}+\mathbf{b}_{4}=\mathbf{w}.$ (21)
@@
Let $\gamma=\mu_{c}\mu_{m}>0$, which is a positive constant independent of $t$ and depends only on hyperparameters $p_{C}$, $p_{M}$, $d_{M}$, and the adaptive parameter learning. Finally, for NN independent offspring:
@@ Appendix I Sensitivity Analysis of Learning Rate
We introduce ABOM-PT, a pre-trained variant of ABOM, where meta-optimization knowledge is distilled from the training set. Specifically, ABOM is executed on 10 source tasks for $T=250$ iterations per task, generating 25002500 prior training samples $\mathcal{M}=\{(P_{k}^{(t)},F_{k}^{(t)})\mid P_{k}^{(t)}\in\mathbb{R}^{N\times d},F_{k}^{(t)}\in\mathbb{R}^{N},k=1,...,K,t=1,...,T\}$. The pre-training objective minimizes the prediction error of population evolution: